2s+3t=10 and 7s-6t=24 elimination method?
Answer by Alan3354(69443)

2s+3t=10 and 7s-6t=24 elimination method?
-------------------
Multiply the 1st eqn by 2, than add the 2nd.
4s + 6t = 20
7s - 6t = 24
--------------- Add
etc